# Rice-like complexity lower bounds for Boolean and uniform automata networks

### v1

@@ 7 Conclusion and perspectives
We have proven Theorem 1, that any non-trivial (in the deterministic setting) or arborescent (in the non-deterministic setting) MSO question on the dynamics of a given automata network (AN) or non-deterministic automata network (NAN) is ${\mathsf{NP}}$-hard or ${\mathsf{coNP}}$-hard to answer, when given the network’s description in the form of a circuit, even in the case of bounded alphabets. Part c of Theorem 1 was proven in [10], on which our developments are grounded. In Section 4 we have included the deterministic constaint to the abstract pumping technique in order to obtain Theorem 1.a. In Section 5 we have treated the arithmetics of the intersection between the sizes of $q$-uniform automata networks, and the sizes of models obtained by pumping. This led to the proof of Theorem 1.b, 1.d in Section 6. In particular, the Rice-like complexity lower bounds also hold for Boolean (deterministic and non-deterministic) automata networks, a model widely used in applications to system biology, for example in the study of gene regulatory networks. Our results intuitively state that any “interesting” computational investigation of the dynamics in a model of interacting entities, is limited to a “high” time-complexity cost in the worst case. Hence there cannot exist any general efficient algorithm, unless some unsuspected complexity collapse occurs (such as ${\mathsf{P}}={\mathsf{NP}}$).
The deterministic case is now closed for MSO (remark that deterministic implies treewidth at most $2$), because the trivial versus non-trivial dichotomy is sharp and deep: all trivial formulas are answers with $\mathcal{O}(1)$ algorithms, whereas all non-trivial formulas are hard for the first level of the polynomial hierarchy (either ${\mathsf{NP}}$ or ${\mathsf{coNP}}$). However, in the non-deterministic case there are computationaly hard non-arborescent formulas (e.g. being a clique), hence the dichotomy does not reaches the sharpness of an “a la Rice” statement. To go further, it would be meaningful to extend the proof technique to other structural graph parameters, such as cliquewidth, rankwidth and twinwidth. In [10], it is formally argued that graph parameterization is necessary to derive general complexity lower bounds for non-deterministic automata networks, because a proof of the basic trivial versus non-trivial dichotomy in this setting implies unexpected complexity collapses (regardless of our considerations on alphabet sizes, and even at first-order).
@@
[…]
@@ Acknowledgments
The authors are thankful to Colin Geniet, Guilhem Gamard and Guillaume Theyssier for helpful and stimulating discussions on technical and non-technical aspects of the present work. They also thank the support of HORIZON-MSCA-2022-SE-01 101131549 ACANCOS and STIC AmSud CAMA 22-STIC-02 (Campus France MEAE) projects.

### v2

@@ 7 Conclusion and perspectives
We have proven Theorem 1, that any non-trivial (in the deterministic setting) or cw-non-trivial (in the non-deterministic setting) MSO question on the dynamics of a given automata network (AN) or non-deterministic automata network (NAN) is $\mathsf{NP}$}-hard or $\mathsf{coNP}$}-hard to answer, when given the network’s description in the form of a circuit, even in the case of bounded alphabets. Part c of Theorem 1 was proven in [10], on which our developments are grounded. In Section 4, we have included the deterministic constraint to the abstract pumping technique in order to obtain Theorem 1.a. In Section 5 we have treated the arithmetics of the intersection between the sizes of $q$-uniform automata networks, and the sizes of models obtained by pumping. This led to the proof of Theorem 1.b,d in Section 6. In particular, the Rice-like complexity lower bounds also hold for Boolean (deterministic and non-deterministic) automata networks, a model widely used in applications to system biology, for example in the study of gene regulatory networks. Our results intuitively state that any “interesting” computational investigation of the dynamics in a model of interacting entities, is limited by a “high” time-complexity cost in the worst case. Hence there cannot exist any general efficient algorithm, unless some unsuspected complexity collapse occurs (such as ${\mathsf{P}}={\mathsf{NP}}$).
The deterministic case is now closed for MSO (remark that deterministic implies cliquewidth at most $6$), because the trivial versus non-trivial dichotomy is sharp and deep: all trivial formulas are answers with $\mathcal{O}(1)$ algorithms, whereas all non-trivial formulas are hard for the first level of the polynomial hierarchy (either $\mathsf{NP}$} or $\mathsf{coNP}$}). However, in the non-deterministic case there are computationally hard cw-trivial formulas (e.g. being a grid), hence the dichotomy does not reach the sharpness of an “a la Rice” statement. To go further, it would be meaningful to extend the proof technique to other structural graph parameters, such as rankwidth and twinwidth. In [10], it is formally argued that graph parameterization is necessary to derive general complexity lower bounds for non-deterministic automata networks, because a proof of the basic trivial versus non-trivial dichotomy in this setting implies unexpected complexity collapses (regardless of our considerations on alphabet sizes, and even at first-order).
@@
Finally, the signature of our logics contains (apart from the $\in$” of MSO) only the two binary relations $\{=,\to\}$ on pairs of configurations (vertices), consequently all expressible questions are valid up to graph isomorphism. Distinguishing configurations may be meaningful for applications, therefore the addition of finer relations is another interesting research track to expand the results.
@@ Acknowledgments
The authors are thankful to Colin Geniet, Guilhem Gamard and Guillaume Theyssier for helpful and stimulating discussions on technical and non-technical aspects of this work. They also thank the support of projects ANR-24-CE48-7504 ALARICE, HORIZON-MSCA-2022-SE-01 101131549 ACANCOS and STIC AmSud CAMA 22-STIC-02.